OpenStudy (anonymous):

3/5(k-2)-1/4(2k-7)<=3. solve and write in interval

OpenStudy (anonymous):

3/5(k-2)-1/4(2k-7)<=3 3/5k-1.2-1/2k-1.75<=3 1/10k-2.95<=3 1/10k<=5.95 \[k \le59.5\] Interval notation \[(59.5,-\infty)\]

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

Hope I helped you!

OpenStudy (anonymous):

why decimals for opening 5(k-2)

OpenStudy (anonymous):

3/5(k-2)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

these is how i was doing it \[[3/(5k-10)]-[1/(4k-28)\le3\] since \[a/b+c/d=(ad+bc)/bd\] \[[3*(8k-28)-(5k-2)]/[5*4(k-2)(2k-7)]\le3\] \[[19k-74]/20(k-2)(2k-7)\le3\] \[(19k-74)/60(k-2)(2k-7)\le1\] critical values \[k=2;(3,5);(3,897)\] \[k<2\] \[7/2<k<74/19\]

OpenStudy (anonymous):

Jonask that is incorrect